A miliammeter of range 10mA has a coil of 1 ohm. To use it as an ammeter of range 1A, the required shunt must have a resistance of?

Grade:12

1 Answers

Khimraj
5 years ago
let shunt resistanceis S then
S(I – Ig) = G*Ig
S*(1 – 0.01) = 1*0.01
S = 1/99 ohm
Hope it clears. If you like answer then please approve it.